rewrite 8x+2y=14 so that y is a function of x

Answer by nerdybill(7384)

Simply isolate 'y' to one side of the equation.
.
Starting with:
8x+2y=14
subtract 8x from both sides:
2y=14-8x
now, divide both sides by 2:
y=7-4x
or
y = -4x + 7